Solve the differential equation y''=1+(y')^2
1
Expert's answer
2020-02-09T16:51:39-0500

We want to solve

"y''=1+(y')^2"


The usual thing is to replace "y'=v" and therefore we get

"v'=1+v^2."


This ODE can be solved by the usual methods, just use what you know. For example separation of variables i.e. solve for v

"\int \frac{dv}{1+v^2} = \int dx"

"arctan(v)=x+C."


We get "v(x)=\tan(x+C)" where C is some constant. Now resubstitute i.e.


"y'=v=\tan(x+C)"


Just integrate both sides with respect to x i.e.

"y(x)=\int \tan(x+C) \ dx"

This can be solved easily, for example use substitution "s=\cos(x+C)"


"\int \tan(x+C) \ dx=-\int \frac{ds}{s}=-\log(s)=-\log(\cos(x+C))"